Solve this system by Substitution.
y = x^2 - 8 and 2x + y = 6

Found 2 solutions by Timnewman, TeachMath:
Answer by Timnewman(323)

Solve this system by Substitution.
y = x^2 - 8 and 2x + y = 6
sol:
y=x^2-8 ---1
2x+y=6 ---2
from 2,
y=6-2x ---3
put 3 in 1
ie, x^2-8=6-2x
x^2+2x-8-6=0
x^2+2x-14=0
then solve the resulting quadratic equation.
GOOD LUCK!

Answer by TeachMath(96)


Solve this system by Substitution.
y = x^2 - 8 and 2x + y = 6

y = x^2 - 8____1
2x + y = 6_____2
Substitute value for y into 2 to get: 2x + x^2 - 8 = 6
x^2 + 2x - 14 = 0
Solve the quadratic equation by using the quadratic equation formula, completing the square, or graphing.